OpenStudy (itiaax):

Given that the sum of the angles A, B and C of a triangle is π radians, show that: sin A + sin B + sin C = sin(A+B) + sin(B+C) + sin(A+C) Please show all working clearly. MEDAL WILL BE AWARDED

OpenStudy (anonymous):

Lets use a 30-60-90 triangle: A = 30 degrees B = 60 degrees C = 90 degrees sin(30 degrees+60 degrees)+sin(60 degrees+90 degrees)+sin(30 degrees+90 degrees) sin(90 degrees)+sin(150 degrees)+sin(120 degrees) sin(360 degrees) = 0 = sin(180 degrees)

OpenStudy (itiaax):

Suppose we don't want to use actual values?

OpenStudy (anonymous):

I suppose you will use sin(A+B) princ. then. sin(A+B) = sin A cos B + cos A sin B

OpenStudy (anonymous):

sin(A+B) = sin A cos B + cos A sin B sin(B+C) = sin B cos C + cos B sin C sin(A+C) = sin A cos C + cos A sin C

OpenStudy (raden):

A+B+C=π therefore, A+B = π-C A+C = π-B B+C = π-A now the righ side can be sin(A+B) + sin(B+C) + sin(A+C) = sin(π-C) + sin(π-B) + sin(π-A) then use the identity : sin(π-x ) = sin(x) so, sin(π-C) + sin(π-B) + sin(π-A) = sin(C) + sin(B) + sin(A) = LHS QED

OpenStudy (itiaax):

What does QED mean?

OpenStudy (raden):

opsss... i mistake in copying and vaste the equation :) that should like this : sin(A+B) + sin(B+C) + sin(A+C) = sin(π-C) + sin(π-A) + sin(π-B) = sin(C) + sin(A) + sin(B)

OpenStudy (raden):

*paste

OpenStudy (raden):

and, actually QED = Quod Erat Demonstrandum it means the proof is complete. :)

OpenStudy (itiaax):

Thanks so much! :)

OpenStudy (raden):

you're welcome